1-If the consumption function is given by C = 150 + 0.85Y and Y increases by 1 unit, then C increases by:

2-Assume that the consumption function is given by C = 150 + 0.85(Y – T) and the tax function is given by T = t0 + t1Y. If t0 increases by 1 unit, then consumption:

3- Assume that the consumption function is given by C = 200 + 0.7(Y – T), the tax function is given by T = 100 + t1Y, and Y = 50K05 L 0.5, where K = 100 and L = 100. If t1 increases from 0.2 to 0.25, then consumption decreases by:

4- Assume that the investment function is given by I = 1,000 – 30r, where r is the real rate of interest (in percent). Assume further that the nominal rate of interest is 10 percent and the inflation rate is 2 percent. According to the investment function, investment will be:
Answer #1

(1) The coefficient of Y in consumption function is the marginal propensity to consume (MPC), which is

MPC = Increase in consumption / Increase in Y

0.85 = Increase in consumption / 1

Increase in consumption = 1 x 0.85 = 0.85 units

(2) C = 150 + 0.85 x [Y - (t0 + t1Y)] = 150 + 0.85 x [(1 - t1)Y - t0] = 150 + 0.85 x (1 - t1)Y - 0.85t0

Therefore if t0 increases by 1 unit, consumption decreases by 0.85 units.

(3) When L = K = 100, Y = 50 x (100)0.5 x (100)0.5 = 50 x 100 = 5,000

C = 200 + 0.7 x [Y - (100 + t1Y)] = 200 + 0.7 x [(1 - t1)Y - 100] = 200 + 0.7 x (1 - t1)Y - 70 = 130 + 0.7 x (1 - t1)Y

= 130 + 0.7 x (1 - t1) x 5,000 = 130 + 3,500 x (1 - t1)

When t1 = 0.2, C = 130 + 3,500 x (1 - 0.2) = 130 + 3,500 x 0.8 = 130 + 2,800 = 2,930

When t1 = 0.25, C = 130 + 3,500 x (1 - 0.25) = 130 + 3,500 x 0.75 = 130 + 2,625 = 2,755

Decrease in consumption = 2,930 - 2,755 = 175

(4) Real interest rate (r) = Nominal interest rate - Inflation rate = 10% - 2% = 8%

Investment (I) = 1,000 - (30 x 8) = 1,000 - 240 = 760